OpenStudy (anonymous):

Simplify. √108 - 3√8 -√27 3√3 - 5√2 -3√3 - 5√2 3√3 - 6√2 -3√3 - 6√2 Can someone help me understand the answer? Im told it is the 3rd one but i didnt get an explanation or work to show me how it is that.

OpenStudy (anonymous):

Minimal polynomial x^4-198 x^2+2025

Parth (parthkohli):

Start with some factoring\[\sqrt{3 \times 2\times 2 \times 3 \times 3} - 3\sqrt{2 \times 2 \times 2} - \sqrt{3 \times 3 \times 3}\]

OpenStudy (anonymous):

i factored and got 6√2 , -6√2 , and 3√3

OpenStudy (anonymous):

^

OpenStudy (anonymous):

but i think the last one is a negative and wouldnt the first 2 cross each other out?

Parth (parthkohli):

Erm well... check the 6√2 again.

OpenStudy (anonymous):

oh i see would it be 6√3 ?

Parth (parthkohli):

Yes! :-)

OpenStudy (anonymous):

Is that your answer 6 sqrt 3?

OpenStudy (anonymous):

oh okay i see now thanks @ParthKohli

Parth (parthkohli):

No, just the first term. And the last term is negative.

Parth (parthkohli):

Here are your terms:\[6\sqrt{3} - 6\sqrt 2 - 3\sqrt{3}\]

Parth (parthkohli):

You can simplify that expression.